# FOF-X: Towards Real-time Detailed Human Reconstruction from a Single Image

### v1

@@ III-C2 FOF to Mesh with Laplacian Coordinate Constraint
However, it is crucial to note that the occupancy field generated by FOF exhibits a view-dependent bias, resulting in observable stair-step artifacts on the reconstructed mesh. This bias marginally impacts the geometric accuracy and only detracts from the visual quality. As illustrated in Fig. 6, when applying Marching Cubes on FOF, only the points on the edges parallel to the $z$-axis remain precise. To address this, we propose a Laplacian coordinate constraint to eliminate the view bias. We adapt the Marching Cubes algorithm to incorporate an additional attribute for each point, which indicates the reliability of the point’s coordinates. The points produced on the edges parallel to the $z$-axis are reliable, and the points produced on the edges orthogonal to the $z$-axis. For those unreliable points, their coordinates are adjusted to confer greater mesh smoothness. This refinement is accomplished by minimizing the mean square of the Laplacian coordinates, expressed as:
@@
where $D$ is a diagonal matrix with $D_{ii}=d_{i}$ on the diagonal, $d_{i}$ is the number of neighbors of the $i^{\textrm{th}}$ point, $X\in\mathbb{R}^{n\times 3}$ is the coordinate matrix to be optimized, and $A$ is the adjacency matrix. $A_{ij}=1$ if points $v_{i}$ and $v_{j}$ are adjacent, otherwise $A_{ij}=0$. $U$ is the index set of unreliable points, which are to be optimized.
@@ IV-B Implementation Details
The reconstruction network of FOF-X takes dual-sided normal maps and the FOF of SMPL as input. The dimensions of the input are $512\times 512\times(3+3+16)$, and the dimensions of the produced feature map are $128\times 128\times 256$. We adopt HRNet-W48 [52] as the image encoder of our system due to its multi-scale perception ability. Since the SMPL mesh only provides a rough shape prior, the high-frequency components corresponding to geometric details are redundant for the input. Empirically, we find that the first 16 terms of the SMPL mesh is sufficient for the shape prior. The decoder progressively upsamples the feature map to $256\times 256$ and $512\times 512$ using bilinear interpolations and residual convolutional blocks, while retaining a channel number of $256$ throughout the process. Finally, a convolutional layer projects the features to a FOF with a resolution of $512\times 512\times 128$. To cope with single view image as input, we use the off-the-shelf image-to-image translation network used in ICON [3] for generating dual-sided normal maps from the single image.
@@
For consistency with previous works [1, 4, 3], we employ a weak perspective camera. The scale of the weak perspective camera is fixed at $1$ for training and evaluation, resulting in an orthogonal projection. But note that we can transform world space to NDC (Normalized Device Coordinates) space to support perspective cameras straightforwardly. We follow the post-processing for hands described in ECON[4], fusing the hands of the parametric model with the reconstructed body. This strategy is not used in our real-time systems.
TABLE II: Quantitative Comparison with existing methods.
CAPE[51] THuman2.1[49] CustomHumans[50]
[…]
PIFu[1] 1.911 2.265 1.461 0.718 3.589 3.792 2.504 0.542 3.635 3.932 2.365 0.526
PIFuHD [11] 3.279 3.393 1.903 0.629 3.963 3.986 2.620 0.497 4.184 4.115 2.432 0.487
PaMIR [12] 1.426 1.629 1.377 0.754 1.549 1.771 1.720 0.762 1.211 1.404 1.350 0.797
ICON [3] 0.967 1.001 0.909 0.807 0.968 0.881 1.338 0.824 0.798 0.755 1.039 0.841
ICONfl [3] 0.817 0.814 0.755 0.832 0.948 0.860 1.254 0.828 0.811 0.761 0.984 0.838
D-IF [33] 0.758 0.714 0.752 0.844 1.210 1.082 1.332 0.770 1.043 0.955 1.098 0.787
ECONif [4] 0.947 0.956 0.928 0.812 1.186 1.224 1.442 0.795 1.015 1.027 1.106 0.807
ECONex[4] 0.887 0.864 0.844 0.819 1.094 1.049 1.301 0.803 1.055 1.040 1.091 0.797
GTA [32] 0.738 0.761 0.808 0.852 1.514 1.524 1.555 0.731 1.382 1.435 1.332 0.736
SiFU [2] 0.676 0.677 0.733 0.863 1.015 1.004 1.271 0.816 0.880 0.900 0.994 0.828
SiTH [19] 1.016 1.030 1.051 0.793 1.218 1.160 1.469 0.773 1.111 1.061 1.216 0.779
FOF-Base [20] 4.583 3.825 2.345 0.500 3.597 3.408 2.489 0.550 4.677 4.084 2.671 0.487
FOF-SMPL [20] 0.765 0.709 0.649 0.845 0.959 0.849 1.120 0.823 0.850 0.759 0.882 0.833
FOF-X(HR32) (Ours) 0.708 0.654 0.634 0.856 0.858 0.768 1.049 0.841 0.731 0.663 0.783 0.854
FOF-X(HR48) (Ours) 0.716 0.672 0.639 0.855 0.835 0.754 1.033 0.844 0.718 0.658 0.771 0.856
@@ IV-C Metrics
3D IoU
While Chamfer distance measures the surface geometric error, we still need a metric to showcase the volumetric geometric error, as many methods apply volumetric loss to supervise their models. However, computing the exact 3D Intersection over Union (IoU) between two meshes is computationally expensive. Therefore, we use an approximate method by uniformly sampling points in the 3D space. To minimize the effects of sampling errors, we employ a stratified sampling strategy. Specifically, the space is divided into a $256\times 256\times 256$ grid. Within each cell of this grid, we randomly sample one point in it. This approach helps to mitigate system errors and provides a more reliable approximation of the 3D IoU.
@@ IV-D Comparison with the State of the Arts
We compare our method with the state-of-the-art approaches, including PIFu [1], PIFuHD [11], PaMIR[12], ICON [3], ECON[4], D-IF[33], GTA[32], SiFU[2], and SiTH[19] on CAPE[51], THuman2.1[49], and CustomHumans[50].
@@
The quantitative results are summarized in Table II. denotes the re-implemented version in ICON[3]. denotes the re-implemented version with the cosine series and the improved inter-conversion algorithms. FOF-X(HR32) and FOF-X(HR48) are FOF-X model with HRNet-W32 and HRNet-W48 respectively. All methods listed, except PIFuHD[11], are trained on THuman2.0[49]. For PIFu[1] and PaMIR[12], similar to ICON[3], we use the re-implemented versions, as the official releases are trained with additional commercial datasets. Since some methods do not provide the training sample lists, we use THuman2.1 instead of THuman2.0 to evaluate all methods. We use $512\times 512\times 512$ resolution for the input of all these methods. We do not compare our method with [14] and [54] because their codes are not released. For simplicity, we also ignore some works that have already been compared with the state-of-the-art methods listed above. We found that some methods are sensitive to the ratio of the human subject’s size in the image. To ensure consistency, we use a ratio of 0.9 for all methods, corresponding to the 1.8m height normalization mentioned in the datasets section. Ground-truth SMPL meshes are used to better demonstrate the capability of the evaluated methods.
As can be seen, our methods show the lowest reconstruction errors compared to other methods on THuman2.1 and CustomHumans datasets. As mentioned in [19], CAPE provides the fitted SMPL+D meshes instead of the original scanned meshes as the ground-truth, leading to mismatches between the image details and the ground-truth meshes. This may cause a potential bias in evaluation. Despite this, our FOF-X can still achieve a comparable metric on CAPE. Notably, our method demonstrates robustness across all benchmarks, achieving better or comparable performance to other computationally heavy methods.
@@
We conduct a qualitative comparison in Fig. 7. As can be seen, SiTH[19] generally produces good results but sometimes introduces artifacts in the limbs and side shapes. ECON[4] suffers from self-occlusions which is the essential limitation of its optimization-based solution; SiFU[2] cannot recover detailed surfaces, although it can produce accurate rough shapes; D-IF[33] has difficulties dealing with limb shapes due to the way it uses the SMPL prior. In contrast, our method is able to reconstruct plausible 3D human models under different body poses and shapes. In terms of surface quality and geometric accuracy, our method outperforms other state-of-the-art approaches. Notably, our method is much more lightweight than others, allowing for real-time reconstruction.
Fig. 8 shows additional comparison on non-frontal view inputs. Generally, it is more challenging to obtain information about the body and limb shape from non-frontal view inputs. In this case, the results of SiTH[19], D-IF[33], and ECON[4] all experience varying degrees of performance degradation. SiFU is still constrained by the resolution issue in producing detailed surfaces. In contrast, our method is more robust and still achieves good results.
@@
Tab. III shows the comparison results in terms of running time. The symbols used in the table are consistent with those in the quantitative comparison. The original system presented in our conference publication, FOF[20], is implemented in Python. Due to limitations in the original mesh-to-FOF algorithm, the SMPL prior is not supported in the FOF real-time system. Since FOF-SMPL requires the execution of three separate scripts, its running time is not reported. In FOF-X(TRT), the TensorRT-accelerated version of our pipeline, we include additional components such as SMPL estimation, image segmentation, and rendering. However, we implemented and optimized the entire process using C++, OpenGL and CUDA, ensuring that all stages are efficiently accelerated. All methods are tested on a single RTX 3090 GPU. It can be seen that our method runs significantly faster than the other existing methods, with a speed improvement of two orders of magnitude.
TABLE III: Running Times Comparison with existing methods.
PIFu[1] PIFuHD [11] PaMIR [12] ICON [3] ICONfl [3] ECONex[4] ECONif [4]
Time (s) 9.98 15.90 36.48 35.00 35.56 108.99 124.55
D-IF [33] GTA [32] SiFU [2] SiTH [19] FOF(Ours)[20] FOF-X (Ours) FOF-X(TRT) (Ours)
Time (s) 61.15 50.29 62.38 83.18 0.04 0.26 0.03
@@ IV-E1 Number of Terms to Use
In the original FOF, we use $312N+1=$ 31 terms of the Fourier series to represent the occupancy field, which has been shown to be sufficient for maintaining geometric accuracy. However, we observe that the absence of high-frequency components can lead to visual artifacts. Therefore, we use $N=128$ terms of the cosine series in FOF-X. We illustrate the impact of varying $N$ on the accuracy of the approximation in Fig. 9. The metrics for different values of $N$ are also presented in Table IV. When $N\leq 16$, noticeable artifacts appear on the meshes. Noting the geometric details in the ear region, although using $N=32$ preserves most geometric information, we prefer $N=128$ in FOF-X to ensure higher visual quality and eliminate potential artifacts.
TABLE IV: Geometric metrics for different number of FOF components.
@@
Figure 9: Results of different number $N$ of coefficients used in FOF-X. We use $N=128$ in our implementation.
IV-E2 Noise Robustness
IV-E3 Dual-sided Normal Maps
@@
IV-E4 Improved Inter-Conversion Algorithm
Fig. 2 shows the effectiveness and robustness of our improved inter-conversion algorithm. As can be seen, our parallelized mesh-to-FOF algorithm demonstrates strong robustness. With the automaton-based discontinuity matcher, it effectively avoids the floating masses caused by duplicate or missing discontinuities. This algorithm is integrated into the real-time pipeline for its excellent parallelizability. Also, our improved FOF-to-mesh algorithm avoids the stair-step artifacts caused by view bias effectively. Although the Laplacian coordinate constraint solution is a little time-consuming, we can use Laplacian smoothing in the real-time pipeline as an alternative and apply our full FOF-to-mesh algorithm during the mesh export stage.

### v2

@@ III-C2 FOF to Mesh with Laplacian Coordinate Constraint
However, it is crucial to note that the occupancy field generated by FOF exhibits a view-dependent bias, resulting in observable stair-step artifacts on the reconstructed mesh. This bias marginally impacts the geometric accuracy and only detracts from the visual quality. As illustrated in Fig. 7, when applying Marching Cubes on FOF, only the points on the edges parallel to the $z$-axis remain precise. To address this, we propose a Laplacian coordinate constraint to eliminate the view bias. We adapt the Marching Cubes algorithm to incorporate an additional attribute for each point, which indicates the reliability of the point’s coordinates. The points produced on the edges parallel to the $z$-axis are reliable, and the points produced on the edges orthogonal to the $z$-axis are not. For those unreliable points, their coordinates are adjusted to confer greater mesh smoothness. This refinement is accomplished by minimizing the mean square of the Laplacian coordinates, expressed as:
@@
where $D$ is a diagonal matrix with $D_{ii}=d_{i}$ on the diagonal, $d_{i}$ is the number of neighbors of the $i^{\textrm{th}}$ point, $X\in\mathbb{R}^{n\times 3}$ is the coordinate matrix to be optimized, and $A$ is the adjacency matrix. $A_{ij}=1$ if points $v_{i}$ and $v_{j}$ are adjacent, otherwise $A_{ij}=0$. $U$ is the index set of unreliable points, which are to be optimized. Unlike excessive Laplacian smoothing iterations that degrade geometric features into spherical approximations, our Laplacian coordinate constraint preserves topological fidelity and guarantees convergence to ground-truth mesh shapes.
@@
TABLE II: Quantitative Comparison with existing methods.
CAPE[46] THuman2.1[47] CustomHumans[48]
[…]
PIFu[1] 2019 9.98 1.911 2.265 1.461 3.589 3.792 2.504 3.635 3.932 2.365
PIFuHD [11] 2020 15.90 3.279 3.393 1.903 3.963 3.986 2.620 4.184 4.115 2.432
PaMIR [12] 2021 36.48 1.426 1.629 1.377 1.549 1.771 1.720 1.211 1.404 1.350
ICON [3] 2022 35.00 0.967 1.001 0.909 0.968 0.881 1.338 0.798 0.755 1.039
ICONfl [3] 2022 35.56 0.817 0.814 0.755 0.948 0.860 1.254 0.811 0.761 0.984
D-IF [30] 2023 61.15 0.758 0.714 0.752 1.210 1.082 1.332 1.043 0.955 1.098
ECONif [4] 2023 124.55 0.947 0.956 0.928 1.186 1.224 1.442 1.015 1.027 1.106
ECONex[4] 2023 108.99 0.887 0.864 0.844 1.094 1.049 1.301 1.055 1.040 1.091
GTA [29] 2023 50.29 0.738 0.761 0.808 1.514 1.524 1.555 1.382 1.435 1.332
SiFU [2] 2024 62.38 0.676 0.677 0.733 1.015 1.004 1.271 0.880 0.900 0.994
SiTH [18] 2024 83.18 1.016 1.030 1.051 1.218 1.160 1.469 1.111 1.061 1.216
PSHuman [43] 2025 76.69 2.208 1.922 0.718 1.464 1.276 0.675 2.066 1.815 0.772
FOF-Base [19] 2022 0.09 4.583 3.825 2.345 3.597 3.408 2.489 4.677 4.084 2.671
FOF-SMPL [19] 2022 - 0.765 0.709 0.649 0.959 0.849 1.120 0.850 0.759 0.882
FOF-X(HR32) (Ours) - 0.02 0.708 0.654 0.634 0.858 0.768 1.049 0.731 0.663 0.783
FOF-X(HR48) (Ours) - 0.02 0.716 0.672 0.639 0.835 0.754 1.033 0.718 0.658 0.771
@@ IV-B Implementation Details
The reconstruction network of FOF-X takes dual-sided normal maps and the FOF of SMPL as input. The dimensions of the input are 512×512×(3+3+16)512$\times$ 512$\times$(3+3+16)512 × 512 × ( 3 + 3 + 16 ), and the dimensions of the produced feature map are 128×128×256128$\times$ 128$\times$ 256128 × 128 × 256. We adopt HRNet-W48 [49] as the image encoder of our system due to its multi-scale perception ability. Since the SMPL mesh only provides a rough shape prior, the high-frequency components corresponding to geometric details are redundant for the input. Empirically, we find that the first 16 terms of the SMPL mesh is sufficient for the shape prior. The decoder progressively upsamples the feature map to $\times 256$ 256 and $\times 512$ 512 using bilinear interpolations and residual convolutional blocks, while retaining a channel number of 256256256 throughout the process. Finally, a convolutional layer projects the features to a FOF with a resolution of 512×512×128512$\times$ 512$\times$ 128512 × 512 × 128. To cope with single view image as input, we use the off-the-shelf image-to-image translation network used in ICON [3] for generating dual-sided normal maps from the single image.
@@
For consistency with previous works [1, 4, 3], we employ a weak perspective camera. The scale of the weak perspective camera is fixed at 111 for training and evaluation, resulting in an orthogonal projection. But note that we can transform world space to NDC (Normalized Device Coordinates) space to support perspective cameras straightforwardly. We follow the post-processing for hands described in ECON[4], fusing the hands of the parametric model with the reconstructed body. This strategy is not used in our real-time systems.
@@ IV-D Comparison with the State of the Arts
We compare our method with the state-of-the-art approaches, including PIFu [1], PIFuHD [11], PaMIR[12], ICON [3], ECON[4], D-IF[30], GTA[29], SiFU[2], SiTH[18], and PSHuman[43] on CAPE[46], THuman2.1[47], and CustomHumans[48].
@@
The quantitative results are summarized in Table II. denotes the re-implemented version in ICON[3]. denotes the re-implemented version with the cosine series and the improved inter-conversion algorithms. FOF-X(HR32) and FOF-X(HR48) are FOF-X model with HRNet-W32 and HRNet-W48 respectively. All methods listed, except PIFuHD[11] and PSHuman[43], are trained on THuman2.0[47]. It is worth noting that PSHuman’s[43] dataset includes THuman 2.1. For PIFu[1] and PaMIR[12], similar to ICON[3], we use the re-implemented versions, as the official releases are trained with additional commercial datasets. Since some methods do not provide the training sample lists, we use THuman2.1 instead of THuman2.0 to evaluate all methods. We use $\times 512\times 512$ 512 resolution for the input of all these methods. We do not compare our method with [13] and [51] because their codes are not released. For simplicity, we also ignore some works that have already been compared with the state-of-the-art methods listed above. We found that some methods are sensitive to the ratio of the human subject’s size in the image. To ensure consistency, we use a ratio of 0.9 for all methods, corresponding to the 1.8m height normalization mentioned in the datasets section. Ground-truth SMPL meshes are used to better demonstrate the capability of the evaluated methods.
As can be seen, our methods show the lowest reconstruction errors compared to other methods on CustomHumans datasets. As mentioned in [18], CAPE provides the fitted SMPL+D meshes instead of the original scanned meshes as the ground-truth, leading to mismatches between the image details and the ground-truth meshes. This may cause a potential bias in evaluation. Despite this, our FOF-X can still achieve a comparable metric on CAPE. In addition, on the THuman2.1 dataset, we are only slightly inferior to PSHuman[43] in terms of normals. However, we note that PSHuman’s[43] training dataset incorporates THuman2.1 without providing clear train/test splits, potentially allowing some test data to be included in training. Notably, our method demonstrates robustness across all benchmarks, achieving better or comparable performance to other computationally heavy methods.
@@
We conduct a qualitative comparison in Fig. 8. As can be seen, SiTH[18] and PSHuman[43] generally produces good results but sometimes introduces artifacts in the limbs and side shapes. ECON[4] suffers from self-occlusions which is the essential limitation of its optimization-based solution; SiFU[2] cannot recover detailed surfaces, although it can produce accurate rough shapes; D-IF[30] has difficulties dealing with limb shapes due to the way it uses the SMPL prior. In contrast, our method is able to reconstruct plausible 3D human models under different body poses and shapes. In terms of surface quality and geometric accuracy, our method outperforms other state-of-the-art approaches. Notably, our method is much more lightweight than others, allowing for real-time reconstruction.
Fig. 9 shows additional comparison on non-frontal view inputs. Generally, it is more challenging to obtain information about the body and limb shape from non-frontal view inputs. In this case, the results of SiTH[18], D-IF[30], and ECON[4] all experience varying degrees of performance degradation. SiFU[2] is still constrained by the resolution issue in producing detailed surfaces. PSHuman[43] has certain difficulties in processing limb shapes. In contrast, our method is more robust and still achieves good results.
@@
Table II shows the comparison results in terms of running time. All methods except PSHuman[43] are tested on a single RTX-3090 GPU, where FOF-X represents the acceleration of FOF-X using TensorRT. PSHuman[43] is tested on the A100 GPU due to memory limitations. Our FOF-X takes 0.02 seconds per frame, demonstrating 500× acceleration over the fastest baseline (PIFu[1]: 9.98s). The original system presented in our conference publication, FOF[19], is implemented in Python. Due to limitations in the original mesh-to-FOF algorithm, the SMPL prior is not supported in the FOF real-time system. Since FOF-SMPL requires the execution of three separate scripts, its running time is not reported. Our enhanced pipeline FOF-X adds additional components such as SMPL estimation, image segmentation, and rendering while maintaining real-time performance. Notably, all variants of our method (Python/TensorRT) consistently outperform existing works in terms of efficiency by two orders of magnitude.
@@ IV-E1 Number of Terms to Use
In the original FOF, we use $2N+1=31$ terms of the Fourier series to represent the occupancy field, which has been shown to be sufficient for maintaining geometric accuracy. However, we observe that the absence of high-frequency components can lead to visual artifacts. Therefore, we use $N=128$ terms of the cosine series in FOF-X. We illustrate the impact of varying $N$ on the accuracy of the approximation in Fig. 10. The metrics for different values of $N$ are also presented in Table III. When $N\leq 16$, noticeable artifacts appear on the meshes. Noting the geometric details in the ear region, although using $N=32$ preserves most geometric information, we prefer $N=128$ in FOF-X to ensure higher visual quality and eliminate potential artifacts.
IV-E2 Sampling Scalability
The frequency-domain resolution of FOF can be adaptively inferred according to the time needs of the system, without the need for retraining. We explored the effect of resolution (number of components) on the results taking $N=128$. Fig. 11 illustrates the effect of changing resolution on the accuracy of the results. Metrics of different $N$ values are also given in Table IV.
IV-E3 Noise Robustness
TABLE III: Geometric metrics for different numbers of FOF components.
@@ IV-E3 Noise Robustness
Figure 10: Results of different number $N$ of coefficients used in FOF-X.
TABLE IV: Geometric metrics for different resolution of FOF.
$Res$ 16 32 64 128 256 512
P2S 3.580 1.764 0.885 0.396 0.139 0.028
Chamfer 3.655 1.746 0.889 0.403 0.146 0.031
Figure 11: Results of different resolution $Res$ used in FOF-X. We use $Res=256$ in our implementation.
IV-E4 Dual-sided Normal Maps
@@
IV-E5 Improved Inter-Conversion Algorithm
Fig. 2 shows the effectiveness and robustness of our improved inter-conversion algorithm. As can be seen, our parallelized mesh-to-FOF algorithm demonstrates strong robustness. The quantitative results for two mesh-to-FOF algorithms are also given in Table VI. With the automaton-based discontinuity matcher, it effectively avoids the floating masses caused by duplicate or missing discontinuities. This algorithm is integrated into the real-time pipeline for its excellent parallelizability. Table VII shows the quantitative results of the FOF-to-mesh methods. To better represent the visual effect of the mesh, we render the mesh as a normal diagram and introduce visual metrics [^2] SSIM (structural similarity), PSNR (peak signal-to-noise ratio) and LPIPS (learned perceptual image patch similarity) measure structural preservation, pixel-level accuracy and human perceptual consistency, respectively. to evaluate results. Our improved FOF-to-mesh algorithm avoids the stair-step artifacts caused by view bias effectively. Although the Laplacian coordinate constraint solution is a little time-consuming, we can use Laplacian smoothing in the real-time pipeline as an alternative and apply our full FOF-to-mesh algorithm during the mesh export stage.
@@ IV-E5 Improved Inter-Conversion Algorithm
IV-F The Real-time Pipeline
To perform real-time monocular human reconstruction, we design a four-stage pipeline. In the first stage, we get an image from the video stream with OpenCV [52], use RVM [53] to get a human mask to remove the background, and predict a SMPL model as prior with 4D-humans[54]. RVM is originally a matting method, and we use a threshold of 0.5 to obtain a binary mask. In the second stage, a CUDA-implemented rasterizer is used to render dual-sided normal maps and a FOF representation for the predicted SMPL. The proposed mesh-to-FOF conversion algorithm is integrated into the rasterizer. In the third stage, the pre-trained dual-sided normal estimator and our reconstruction network are encapsulated into a single neural network module, which is used to reconstruct the FOF. In the final stage, the FOF is resized to a proper resolution ($\times 256$ 256 in our implementation) and transformed into a mesh. All four stages are implemented in PyTorch and accelerated with TensorRT, running on a single RTX-4090 GPU. Our pipeline works at 30+ FPS, which is limited by the camera (Logic C920 Pro). With a recorded video, it can achieve a frame rate of about 50 FPS. Fig. 13 shows the real-time reconstruction system (all human subjects provided informed consent for data collection). Please check the demo video for more results.
TABLE VI: Geometric Metrics for Different Mesh-to-FOF algorithms
[…]
Algorithms Used in FOF 0.447 0.206 0.264
Algorithms Used in FOF-X 0.041 0.063 0.099
TABLE VII: Visual Metrics for Different FOF-to-mesh algorithms
[…]
w/o Laplacian coordinate constraint 0.959 32.09 0.018
w/ Laplacian coordinate constraint 0.965 32.35 0.015